# Monte Carlo Simulation of Operator Dynamics and Entanglement in Dual-Unitary Circuits

### v2

@@ I Introduction
When focusing solely on the statistical properties of operator dynamics, i.e. how the size and support of an operator evolves on average without concern for the specific operator content, more efficient approaches become available [3, 28, 4, 7, 29]. For locally scrambled quantum dynamics [15, 16], where the unitary evolution operators are weakly symmetric [30, 31] under local basis transformations, the operator dynamics can be mapped to an effective classical diffusion model in the space of operator supports [3, 29, 32]. This mapping simplifies the quantum evolution into a tractable classical Markov process, allowing for efficient simulation via Monte Carlo (MC) sampling.
@@
We demonstrate, via Monte Carlo simulations of the local operator density and entanglement in a dual-unitary circuit, that the operator density near the light cone converges exponentially to $\frac{3}{4}$, with a rate determined by the gate parameter $\alpha$. In the long-time limit, we derive analytical bounds for the density profile, validated by Monte Carlo simulations. For operator entanglement, we observe volume-law scaling across different subregions, with the volume-law coefficient’s dependence on $\alpha$ varying. A key finding is the transition from maximal to sub-maximal entanglement growth for left movers at the left light cone, consistent with prior theoretical work [39, 36, 35, 40]. Our study provides a direct mapping of the volume-law coefficient to $\alpha$, confirmed by mean-field (MF) theory and Monte Carlo simulations, offering new insights into operator dynamics and entanglement transitions in dual-unitary circuits.
@@ II.3 Monte Carlo Simulation
assuming the last time step is an odd step without lost of generality. In this way, averaging $I(b^{\prime(1)},b^{\prime(2)})$ also gives unbiased estimation of the operator entanglement $\overline{S_{A}^{(2)}}=-\ln\langle I(b^{\prime(1)},b^{\prime(2)})\rangle$. Since the summation of configurations $b$ can be factorized to that on each pair of sites at $(2k-1,2k)$, which can computed efficiently, we can now obtain a finite real number $I(b^{\prime(1)},b^{\prime(2)})$ from each Monte Carlo sampling instead of waiting an exponentially long time for $\delta_{b^{(1)}_{x}b^{(2)}_{x}}=1$ to occur. Similar approach in spirit, to overcome the exponential observable with log-normal distribution to a typical observable with normal distribution [43, 44, 45], has been employed in the recent computation of entanglement entropy and other exponential observables in the quantum many-body systems within quantum Monte Carlo and Tensor Network setup  [46, 47, 48, 49, 50, 51, 52, 53, 44].
@@ Acknowledgements.
Acknowledgments —  We thank Yimu Bao, Tian-Gang Zhou for discussions. MHS, TTW and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. 17301721, AoE/P-701/20, 17309822, HKU C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22), the GD-NSF (No. 2022A1515011007). PZ acknowledges the support from the National Natural Science Foundation of China under Grant No. 12374477. YZY is supported by a startup fund from UCSD. We thank HPC2021 system under the Information Technology Services, The University of Hong Kong, as well as the Beijing PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported within this paper. ZYZ, YZY and ZYM acknowledge the hospitality of Kavli Institute for Theoretical Physics (KITP) and thank the organizers of the KITP program “Correlated Gapless Quantum Matter” where this work was initiated. This research was supported in part by grant NSF PHY-2309135 to the KITP.

### v3

@@ I Introduction
When focusing solely on the statistical properties of operator dynamics, i.e. how the size and support of an operator evolves on average without concern for the specific operator content, more efficient approaches become available [3, 28, 4, 7, 29, 30]. For locally scrambled quantum dynamics [15, 16], where the unitary evolution operators are weakly symmetric [31, 32] under local basis transformations, the operator dynamics can be mapped to an effective classical diffusion model in the space of operator supports [3, 29, 33]. This mapping simplifies the quantum evolution into a tractable classical Markov process, allowing for efficient simulation via Monte Carlo (MC) sampling.
@@
We demonstrate, via Monte Carlo simulations of the local operator density and entanglement in a dual-unitary circuit, that the operator density near the light cone converges exponentially to $\frac{3}{4}$, with a rate determined by the gate parameter $\alpha$. In the long-time limit, we derive analytical bounds for the density profile, validated by Monte Carlo simulations. For operator entanglement, we observe volume-law scaling across different subregions, with the volume-law coefficient’s dependence on $\alpha$ varying. A key finding is the transition from maximal to sub-maximal entanglement growth for left movers at the left light cone, consistent with prior theoretical work [40, 37, 36, 41]. We provide a concrete physical explanation for the emergence of the entanglement transition based on operator emission, and clarify the setup for observing the transition in the physical regime by focusing on left-movers only. Furthermore, our study provides a direct mapping of the volume-law coefficient to $\alpha$, confirmed by mean-field (MF) theory and Monte Carlo simulations. Our results offering new insights into operator dynamics and entanglement transitions in dual-unitary circuits.
@@ II.3 Monte Carlo Simulation
assuming the last time step is an odd step without lost of generality. In this way, averaging $I(b^{\prime(1)},b^{\prime(2)})$ also gives unbiased estimation of the operator entanglement $\overline{S_{A}^{(2)}}=-\ln\langle I(b^{\prime(1)},b^{\prime(2)})\rangle$. Since the summation of configurations $b$ can be factorized to that on each pair of sites at $(2k-1,2k)$, which can computed efficiently, we can now obtain a finite real number $I(b^{\prime(1)},b^{\prime(2)})$ from each Monte Carlo sampling instead of waiting an exponentially long time for $\delta_{b^{(1)}_{x}b^{(2)}_{x}}=1$ to occur. Similar approach in spirit, to overcome the exponential observable with log-normal distribution to a typical observable with normal distribution [44, 45, 46], has been employed in the recent computation of entanglement entropy and other exponential observables in the quantum many-body systems within quantum Monte Carlo and Tensor Network setup  [47, 48, 49, 50, 51, 52, 53, 54, 45, 55, 56]. All code used for generating the results in this work can be accessed online [57].
@@ Acknowledgements.
Acknowledgments —  We thank Yimu Bao, Tian-Gang Zhou for discussions. MHS, TTW and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. 17301721, AoE/P-701/20, 17309822, HKU C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22), the GD-NSF (No. 2022A1515011007). PZ is supported by the Shanghai Rising-Star Program under grant number 24QA2700300, the NSFC under grant 12374477, and the Innovation Program for Quantum Science and Technology ZD0220240101. YZY is supported by a startup fund from UCSD. We thank HPC2021 system under the Information Technology Services, The University of Hong Kong, as well as the Beijing PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported within this paper. ZYZ, YZY and ZYM acknowledge the hospitality of Kavli Institute for Theoretical Physics (KITP) and thank the organizers of the KITP program “Correlated Gapless Quantum Matter” where this work was initiated. This research was supported in part by grant NSF PHY-2309135 to the KITP.